WEAK C-H⋯ π HYDROGEN BOND AND FREE INTERNAL ROTATION IN THE BENZENE-TRIFLUOROMETANE DIMER

JOSÉ L. ALONSO, JUAN C. LÓPEZ, Grupo de Espectroscopía Molecular (GEM), Departamento de Química Física y Química Inorgánica, Facultad de Ciencias, Universidad de Valladolid, E-47005 Valladolid, Spain; \underline{WALTHER CAMINATI}, Dipartimento di Chimica ”G. Ciamician” dell’Università, Via Selmi 2, I-40126 Bologna, Italy.

The pure rotational spectra of the normal and benzene-^{13}C species of the benzene-trifluoromethane complex have been measured using molecular beam Fourier transform microwave spectroscopy. The normal species is a symmetric top, with the symmetry axes, C_{3} of CHF_{3} and C_{6} of benzene, lying along the same line. The rotational spectrum of the benzene-13C isotopomer is that of a slightly asymmetric top. Both species are characterized by an almost free rotation of the two subunits with respect to each other. From the present data the calculated length of the C-H⋯ π hydrogen bond is 2.366(2) Å. The analysis of the structural and energetic features of the C-H⋯ π interaction allow us to classify it as a weak hydrogen bond.